sqrt(50m^3n^5)
sqrt(72x^3y^4z^5
sqrt(2/7 * 7/30

Answer by checkley75(3666)

sqrt(50m^3n^5)=5mn^2sqrt(2mn)
sqrt(72x^3y^4z^5=6xy^2z^2sqrt(2xz)
sqrt(2/7*7/30)=sqrt(14/210)=sqrt(1/15)